4. Refer to the previous problem. How large a sample would be required to estimate the population proportion to within .05 with 95 percent confidence (.30 is the best available estimate of p): (a) If the finite population correction can be ignored? (b) If the finite population correction is not ignored and N 1500


1
Expert's answer
2021-07-27T06:05:27-0400

The margin of error for constructing a 100(1-α) % confidence interval for population proportion when the finite population correction can be ignored is

"E = Z_{α/2}\times \sqrt{\frac{\hat{p}(1- \hat{p})}{n}}"

"\hat{p}" = sample proportion

n = sample size

α = 1 – conference level

When the finite population correction is not ignored, then the margin of error is

"E = Z_{α/2}\times \sqrt{\frac{\hat{p}(1- \hat{p})}{n}} \times \sqrt{\frac{N-n}{N-1}}"

N= the population size

"α=1-0.95=0.05 \\

Z_{α/2}=Z_{0.05/2}=Z_{0.025}=1.96 \\

E=0.05 \\

\hat{p}=0.30"

(a)

"E = Z_{α/2}\times \sqrt{\frac{\hat{p}(1- \hat{p})}{n}} \\

0.05 = 1.96 \times \sqrt{\frac{0.30(1- 0.30)}{n}} \\

n = \frac{(1.96)^2 \times 0.30 \times 0.70}{(0.05)^2} \\

n=322.69 ≈323"

The required sample size is n=323.

(b) N=1500

"E = Z_{α/2}\times \sqrt{\frac{\hat{p}(1- \hat{p})}{n}} \times \sqrt{\frac{N-n}{N-1}} \\

0.05 = 1.96\times \sqrt{\frac{0.30(1- 0.30)}{n}} \times \sqrt{\frac{1500-n}{1500-1}} \\

(0.05)^2 = (1.96)^2 \times \frac{0.30 \times 0.70}{n} \times \frac{1500-n}{1499} \\

\frac{1500-n}{n}= \frac{(0.05)^2 \times 1499}{(1.96)^2 \times 0.30 \times 0.70} \\

 \frac{1500-n}{n}= 4.64526 \\

\frac{1500}{n}-1= 4.64526 \\

n = \frac{1500}{4.64526+1} \\

n = 265.71 ≈266"

The required sample size is n=266.